### v2

# Time-Aware One Step Diffusion Network for Real-World Image Super-Resolution

Abstract
Diffusion-based real-world image super-resolution (Real-ISR) methods have demonstrated impressive performance. To achieve efficient Real-ISR, many works employ Variational Score Distillation (VSD) to distill pre-trained stable-diffusion (SD) model for one-step SR with a fixed timestep. However, due to the different noise injection timesteps, the SD will perform different generative priors. Therefore, a fixed timestep is difficult for these methods to fully leverage the generative priors in SD, leading to suboptimal performance. To address this, we propose a Time-Aware one-step Diffusion Network for Real-ISR (TADSR). We first introduce a Time-Aware VAE Encoder, which projects the same image into different latent features based on timesteps. Through joint dynamic variation of timesteps and latent features, the student model can better align with the input pattern distribution of the pre-trained SD, thereby enabling more effective utilization of SD’s generative capabilities. To better activate the generative prior of SD at different timesteps, we propose a Time-Aware VSD loss that bridges the timesteps of the student model and those of the teacher model, thereby producing more consistent generative prior guidance conditioned on timesteps. Additionally, though utilizing the generative prior in SD at different timesteps, our method can naturally achieve controllable trade-offs between fidelity and realism by changing the timestep condition. Experimental results demonstrate that our method achieves both state-of-the-art performance and controllable SR results with only a single step.
Introduction
Image Super-Resolution (ISR) aims to reconstruct a high-quality (HQ) image from its low-quality (LQ) counterpart. Different from the simple degradation assumptions in traditional ISR (Lim et al. 2017; Zhang et al. 2018b; Chen et al. 2021; Liang et al. 2021; Chen et al. 2023; Dong et al. 2014), Real-World Image Super-Resolution (Real-ISR) aims to restore HQ images from LQ inputs degraded by complex and unknown factors in real-world scenarios, which has recently attracted increasing attention (Wang et al. 2021; Xie et al. 2023; Zhang et al. 2021; Liang, Zeng, and Zhang 2022).
To address a broader spectrum of degradation types and achieve more realistic results, many researchers have turned to generative models. Recently, diffusion models (Ho, Jain, and Abbeel 2020) have demonstrated a superior ability to generate fine-grained details in image generation tasks. Consequently, several works have explored leveraging the generative priors in pre-trained Stable Diffusion (SD) models (Rombach et al. 2022) to tackle Real-ISR, yielding impressive results (Wang et al. 2023a; Lin et al. 2023; Tao et al. 2023; Wu et al. 2024c; Duan et al. 2025). Nevertheless, the iterative denoising process inherent in diffusion models introduces significant computational overhead and latency.
Figure 1: (a) Comparison between our TADSR and PisaSR. In PisaSR, increasing the semantic weight $\lambda_{sem}$ leads to the restoration of more realistic images. As the timestep condition $t_{s}$ increases, our model demonstrates a significant improvement in generative capability, recovering a more realistic parrot image. In contrast, PisaSR shows only an increase in sharpness as $\lambda_{sem}$ increases. (b) The input image and the corresponding outputs of the teacher model at different timesteps $t$ with the VSD loss. The outputs of the teacher model vary significantly across different timesteps, reflecting distinct guidance orientations.
To overcome these limitations, some researchers have focused on distilling SD into an efficient one-step model for Real-ISR (Wu et al. 2024b; Sun et al. 2025; Chen et al. 2025; Dong et al. 2025; Zhang et al. 2024). Specifically, OSEDiff (Wu et al. 2024b) first leverages the Variational Score Distillation (VSD) loss (Wang et al. 2023b) to distill the SD model, enabling realistic image reconstruction in a single step. Subsequently, S3Diff (Zhang et al. 2024), PisaSR (Sun et al. 2025), AdcSR (Chen et al. 2025), and TSDSR (Dong et al. 2025) also adopt distillation-based approaches to develop SD-based one-step Real-ISR models, using either adversarial loss or modified VSD loss.
However, these works generally fix the injected timestep, e.g. step 999, and use the original VAE encoder in SD, which prevent them from effectively leveraging the generative prior in SD. Specifically, as shown in Fig. 1(b), the output of the teacher model (pre-trained SD) varies depending on the input timestep and latent features. When timestep $t$ equals 100, there are only differences in texture details between the output of the teacher model and the input image. In contrast, when $t$ increases to 300, the output shows clear differences, reflecting more of the leaf-related semantic prior learned by SD. However, with $t$ increasing further to 600, most of the image information is lost, and the teacher model can only recover the overall structure and color of the leaves, failing to provide meaningful guidance. These observations imply that fixed timestep and latent features are insufficient to fully activate the generative priors across different timesteps in SD, the gradient guidance from the teacher model is also timestep-dependent. As a result, as shown in Fig. 1(a), although we increase the semantic weight $\lambda_{sem}$ in PisaSR (Sun et al. 2025), it only produces enhanced sharpness without significantly enriching the semantic content.
In this paper, we propose Time-Aware One Step Diffusion Network for Super-Resolution (TADSR), a framework that more effectively distills the generative prior of SD at different timesteps into a one-step diffusion model for Real-ISR. Our TADSR consists of two key components: (1) Time-Aware VAE Encoder (TAE): we introduce a time embedding layer into the VAE encoder, enabling it to map the same image to different latent representations based on the timestep, thereby achieving coordinated adjustment between the timestep and latent representation. Through the TAE, we can establish a matching relationship between the input timestep and latent representation similar to that in the original SD, allowing for more effective utilization of the generative priors of SD at different timesteps. (2) Time-Aware Variational Score Distillation (TAVSD) Loss: we design a mapping function to associate the timestep injected into the SR network with the one used in the VSD loss. When the SR network is conditioned on a larger timestep, the teacher receives a latent image corrupted with stronger noise, providing guidance that emphasizes stronger semantic generation in the reconstruction results. Conversely, smaller timesteps lead to similar results with reconstruction, primarily enhancing texture details. Therefore, TAVSD can provide more consistent generative guidance condition on the injected timestep in the SR network.
Thanks to TAE and TAVSD, the proposed TADSR can naturally leverage the generative priors of SD at different timesteps to achieve controllable trade-offs between fidelity and realism in Real-ISR and superior performance. As shown in Fig. 1(a), our method gradually generates a more realistic parrot as the timestep increases, though fully leveraging the generative priors in SD.
Related Work
Real-World Image Super-Resolution
Traditional ISR methods (Lim et al. 2017; Zhang et al. 2018b; Chen et al. 2021; Liang et al. 2021; Chen et al. 2023; Dong et al. 2014; Zhao et al. 2025) typically degrade HQ images using simple downsampling operations to construct HQ-LQ image pairs for training, supervised by pixel-level losses. However, these approaches struggle to handle images degraded by complex real-world processes and often result in overly smooth reconstructions. To better simulate the unknown and complex degradations in real-world scenarios, several studies (Zhang et al. 2021; Wang et al. 2021) have proposed more sophisticated degradation pipelines to synthesize LQ data. Specifically, BSRGAN (Zhang et al. 2021) introduces a random combination of basic degradation operations (e.g., downsampling, blurring, noise) injection, with varying intensities to generate more realistic HQ-LQ pairs. Real-ESRGAN (Wang et al. 2021) proposes a second-order degradation scheme to cover a broader range of degradation types. In addition, inspired by Generative Adversarial Networks (GANs), researchers have incorporated discriminators into Real-ISR frameworks and adopted adversarial losses to encourage the reconstruction of more realistic images. Although these GAN-based methods can produce richer texture details compared to traditional approaches, they are often unstable to train and prone to generating unnatural artifacts (Wu et al. 2024b).
Figure 2: Overview of TADSR. We train a Student Model $G_{\theta}$ to perform one-step Real-ISR, which consists of a Time-Aware VAE Encoder $E_{\theta}$ and a UNet $F_{\theta}$. We randomly sample a timestep $t_{s}$ and map it to $t_{v}$. The $t_{s}$ and the LQ image are fed into the encoder $E_{\theta}$ to obtain the LQ latent. Then, $t_{s}$ and the LQ latent are fed into the UNet $F_{\theta}$ to produce the reconstructed latent feature $\hat{z}_{0}$. After adding noise to $\hat{z}_{0}$ corresponding to $t_{v}$, we feed it and $t_{v}$ into the teacher model and the LoRA model to compute the TAVSD loss (orange flow). The reconstruction loss (blue flow) in pixel space and TAVSD loss is then used to jointly update the student model $G_{\theta}$. For the LoRA Model, we employ the diffusion loss (green flow) for training.
Diffusion-Based Real-ISR
Recently, many researchers have leveraged the powerful generative priors of pre-trained diffusion models for Real-ISR tasks to achieve high-fidelity image reconstruction. For example, StableSR (Wang et al. 2023a) conditions the diffusion process on LQ images by injecting them through a learnable time-aware encoder into the SD model, enabling strong detail generation capabilities. DiffBIR (Lin et al. 2023) utilizes ControlNet to extract structural information from LQ images and incorporates it as a control signal to better guide the generative prior of SD for super-resolution. PASD (Tao et al. 2023) and SeeSR (Wu et al. 2024c) extract semantic information from LQ inputs and inject it into SD, resulting in more realistic super-resolved outputs.
Although these approaches yield impressive results, the multi-step denoising process leads to high computational and time costs. To accelerate diffusion-based Real-ISR, OSEDiff introduces the VSD loss to distill the pre-trained SD model, enabling realistic image reconstruction in a single step by only training LoRA parameters mounted on SD. S3Diff further adopts degradation-guided LoRA adapters combined with adversarial training to achieve one-step super-resolution. PisaSR trains two separate LoRA adapters for pixel-level and semantic-level guidance, allowing controllable trade-offs between realism and fidelity.
However, these methods overlook the varying generative capabilities of SD in different timesteps: larger timesteps favor semantic and structural generation, while smaller timesteps emphasize texture and detail. Our work aims to fully exploit these time-dependent generation characteristics to achieve superior SR performance and a natural balance between visual fidelity and generative realism.
Methodology
Problem Definition
Real-ISR aims to reconstruct HQ images $x_{H}$ from LQ images $x_{L}$ that suffer from complex and unknown degradations. With the advancement of deep learning, researchers have commonly adopted neural networks $G_{\theta}$ to estimate the HQ images and optimize the network through loss functions. The general form of the loss function is as:
$\begin{gathered}\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(x_{L},x_{H})\sim\mathcal{D}}[\mathcal{L}_{Rec}(G_{\theta}(x_{L}),x_{H})\\ +\lambda\mathcal{L}_{Reg}(G_{\theta}(x_{L}))],\end{gathered}$ (1)
where the $\mathcal{L}_{Rec}$ denotes the reconstruction loss to optimize the fidelity of the reconstructed results. $\mathcal{L}_{Reg}$ is the regression loss to enhance the realism of the results, and $\lambda$ is a hyperparameter to balance $\mathcal{L}_{Rec}$ and $\mathcal{L}_{Reg}$.
Recently, with the advancement of diffusion models, several studies (Wu et al. 2024b; Sun et al. 2025) have leveraged the generative prior in pre-trained SD and adopted VSD as a regression objective. The formation of VSD is:
$\nabla_{\theta}\mathcal{L}_{VSD}(\hat{z},c)=\mathbb{E}_{t,\epsilon}\left[\omega(t)\left(\epsilon_{\psi}(\hat{z}_{t};t,c)-\epsilon_{\phi}(\hat{z}_{t};t,c)\right)\frac{\partial\hat{z}}{\partial\theta}\right],$ (2)
where $\epsilon_{\psi}$ is the pre-trained diffusion model (teacher model), $\epsilon_{\phi}$ represents its replica with trainable LoRA (LoRA model), and $c$ is a text embedding of a caption describing the input image. $\hat{z}=G_{\theta}(x_{L})$ is the output of the student network $G_{\theta}$, and $\hat{z}_{t}=\alpha_{t}\hat{z}+\beta_{t}\epsilon$ is the noisy latent input. $\epsilon$ is the gaussian noise, and $\alpha_{t}$ and $\beta_{t}$ are the scale parameters in diffusion.
Formally, the VSD loss can be viewed as the residual between the noise outputs of the teacher model and the LoRA model, which is equivalent to the residual of their predicted latent images (detailed proof can be found in supplementary material). Therefore, in the following, we represent the VSD loss using the latent images and their residual.
Overview
As illustrated in Fig. 2, we distill a Student Model $G_{\theta}$ to perform one-step Real-ISR, which is consist of a Time-Aware Encoder $E_{\theta}$ and a UNet $F_{\theta}$. First, we sample an HQ-LQ image pair from the dataset and a timestep $t_{s}$. Then, both the LQ image and $t_{s}$ are fed into Student Model $G_{\theta}$ to obtain the latent output $\hat{z}_{0}$. We decode $\hat{z}_{0}$ into pixel space and compute the reconstruction loss with the HQ image to ensure fidelity. In the latent space, we map the $t_{s}$ to another timestep $t_{v}$, and add noise corresponding to timestep $t_{v}$ to the $\hat{z}_{0}$ to obtain $\hat{z}_{t_{v}}$. Then, we feed $\hat{z}_{t_{v}}$ and $t_{v}$ into both the teacher model and the LoRA model to compute the Time-Aware Variational Score Distillation loss $\mathcal{L}_{TAVSD}$ to enhance the realism. Consistent with Fig. 1(b), when $t_{v}$ is small, the gradients produced by the TAVSD loss are relatively small and mainly reflect texture details. In contrast, when $t_{v}$ is large, the gradients become significantly large and provide more semantic guidance.
Figure 3: PCA visualization of latent features produced by TAE under different timesteps $t_{s}$, and the corresponding mean and standard deviation (Std) of latent features. TAE can encode the same image into distinct latent features conditioned on different timesteps, which aligns with the synchronized variation between timesteps and latent features in the pre-trained SD.
Time-Aware VAE Encoder
In the original SD, as timestep increases, the latent features fed into the model are injected with more noise, thereby activating different generative priors to produce the image. However, current one-step SD-based Real-ISR methods typically adopt a fixed timestep during training, making it difficult to fully exploit the generative priors in SD. Since the original VAE encoder can only encode the same image into a single latent feature, simply sampling the timestep randomly is insufficient, which is clearly inconsistent with the correlation between timesteps and latent distributions in SD.
To better utilize the generative priors in SD, we propose a Time-Aware VAE Encoder (TAE). By incorporating a temporal embedding layer into the VAE encoder, TAE encodes the input image into different latent distributions conditioned on the timestep $t_{s}$, enabling synchronized variation between $t_{s}$ and latent distribution, thus better activating the generative priors at different timesteps within SD. This process can be formulated as:
$z_{L}=E_{\theta}(x_{L},t_{s}),\hat{z}=F_{\theta}(z_{L},t_{s}),$ (3)
where $E_{\theta}$ is the TAE model and $F_{\theta}$ is the Unet model.
Figure 4: (a) Mean and standard deviation (Std) of the VSD loss at different timesteps. (b) The outputs of the teacher model and the LoRA model are decoded into pixel space and gradients in latent space at different timesteps $t$.
As shown in Fig. 3, with the same input image, TAE encodes it into different latent feature condition on the timestep $t_{s}$. Overall, as the $t_{s}$ increases, both the mean and variance of the latent features show a decreasing trend. After visualizing the latent feature via PCA dimensionality reduction, we can also clearly observe the changes in the latent space.
Time-Aware Variational Score Distillation
Following the OSEDiff (Wu et al. 2024b), the VSD loss has been widely adopted in SD-based one-step Real-ISR methods to enhance the realism of reconstruction results. However, these methods generally overlook that the guidance provided by VSD at different timesteps is actually inconsistent. As shown in Fig. 4, the mean and standard deviation of the VSD loss exhibit a clear upward trend as the timestep increases. When decoding the results at different timesteps into the pixel space, we observe that at $t_{v}$ equals 100, the outputs of the teacher and LoRA models are similar, and the gradients mainly reflect enhancements in texture details. In contrast, when $t_{v}$ increases to 300, the teacher model’s output contains significantly more semantic information while the LoRA model’s output remains smooth, and the gradients reflect global semantic guidance. However, when $t_{v}$ increases to 600, the teacher model can only recover coarse color and structural information from the noisy latent input, making it difficult to provide meaningful guidance. This implies that the VSD loss provides distinct guidance for the same image depending on $t_{v}$. Such opposing directional guidance creates conflicting optimization signals for the student model, ultimately leading to suboptimal convergence.
Table 1: A comprehensive evaluation against state-of-the-art methods across synthetic and real-world datasets. The top-performing and runner-up results under each metric are marked in red and blue, respectively.
Datasets Metrics StableSR DiffBIR SeeSR SinSR S3Diff OSEDiff PisaSR TSDSR AdcSR TADSR
DIV2k-Val
PSNR $\uparrow$ 23.261 23.409 23.679 24.417 23.530 23.723 23.867 22.17 23.743 23.815
SSIM $\uparrow$ 0.5726 0.5732 0.6043 0.6023 0.5933 0.6109 0.6058 0.5602 0.6017 0.6028
LPIPS $\downarrow$ 0.3113 0.3456 0.3194 0.3235 0.2581 0.2942 0.2823 0.2736 0.2853 0.3078
CLIPIQA $\uparrow$ 0.6771 0.7082 0.6935 0.6505 0.7001 0.6682 0.6928 0.7149 0.6763 0.7353
MUSIQ $\uparrow$ 65.918 68.396 68.665 62.838 67.923 67.971 69.681 70.65 67.995 69.649
MANIQA $\uparrow$ 0.6174 0.6355 0.6222 0.5392 0.6311 0.6132 0.6375 0.6077 0.6073 0.6443
TOPIQ $\uparrow$ 0.5979 0.6344 0.6856 0.5721 0.6334 0.6188 0.6619 0.6672 0.6526 0.7044
QALIGN $\uparrow$ 3.5273 3.8774 3.9765 3.5159 3.8666 3.8357 3.8812 3.927 3.612 4.0783
DRealSR
PSNR $\uparrow$ 28.030 25.929 28.073 28.345 27.539 27.915 28.318 26.20 28.099 28.387
SSIM $\uparrow$ 0.7536 0.6526 0.7684 0.7491 0.7491 0.7833 0.7804 0.7170 0.7726 0.7758
LPIPS $\downarrow$ 0.3284 0.4518 0.3173 0.3697 0.3109 0.2968 0.2960 0.3116 0.3046 0.3235
CLIPIQA $\uparrow$ 0.6356 0.6863 0.6909 0.6375 0.7131 0.6974 0.6971 0.7309 0.7049 0.7398
MUSIQ $\uparrow$ 58.512 65.667 65.080 55.009 63.941 64.691 66.113 66.12 66.266 67.016
MANIQA $\uparrow$ 0.5594 0.6289 0.6051 0.4894 0.6124 0.5903 0.6160 0.5820 0.5915 0.6309
TOPIQ $\uparrow$ 0.5323 0.6149 0.6575 0.5122 0.6040 0.6002 0.6333 0.6251 0.6527 0.6758
QALIGN $\uparrow$ 3.0614 3.6011 3.5882 3.0982 3.6148 3.5450 3.5838 3.6928 3.6520 3.7491
RealSR
PSNR $\uparrow$ 24.645 24.240 25.149 26.266 25.183 25.148 25.503 23.404 25.469 25.166
SSIM $\uparrow$ 0.7080 0.6649 0.7211 0.7341 0.7269 0.7338 0.7418 6886 0.7301 0.7150
LPIPS $\downarrow$ 0.3002 0.3470 0.3007 0.3241 0.2721 0.2920 0.2672 0.2805 0.2885 0.3168
CLIPIQA $\uparrow$ 0.6234 0.6959 0.6699 0.6153 0.6731 0.6687 0.6699 0.7199 0.6731 0.7283
MUSIQ $\uparrow$ 65.883 68.340 69.819 60.575 65.824 69.087 70.147 70.7710 69.899 71.182
MANIQA $\uparrow$ 0.6230 0.6530 0.6450 0.5409 0.6427 0.6337 0.6551 0.6311 0.6353 0.6715
TOPIQ $\uparrow$ 0.5748 0.6052 0.6890 0.5188 0.6162 0.6251 0.6374 0.6642 0.6793 0.7082
QALIGN $\uparrow$ 3.2767 3.6313 3.7190 3.1889 3.6638 3.6915 3.6355 3.7748 3.7749 3.9477
RealLR200
CLIPIQA $\uparrow$ 0.6036 0.7072 0.7023 0.6474 0.7122 0.6792 0.7153 0.7248 0.7048 0.7741
MUSIQ $\uparrow$ 62.863 67.727 70.195 63.126 68.897 69.041 70.935 70.930 69.759 72.166
MANIQA $\uparrow$ 0.5922 0.6464 0.6482 0.5522 0.6536 0.6331 0.6639 0.6363 0.6354 0.6738
TOPIQ $\uparrow$ 0.5286 0.5905 0.6900 0.5689 0.6401 0.5990 0.6627 0.6664 0.6684 0.7249
QALIGN $\uparrow$ 3.3409 3.7782 4.0305 3.4105 3.9228 3.8459 3.9891 3.8908 3.9312 4.2622
This phenomenon arises because most of the image information is preserved when $t$ is small, and both the teacher and LoRA models focus mainly on generating texture details. As $t$ increases, the injection of more noise gradually obscures the underlying image content, forcing the teacher model to rely more heavily on its generative prior. However, since the LoRA model is trained on low-quality data generated by the student model and does not employ the CFG strategy (Ho and Salimans 2022), its outputs tend to be overly smooth.
Considering that the guidance provided by VSD varies across different timesteps, we establish a connection between the timestep $t_{s}$ input to the student model and the timestep $t_{v}$ in the teacher model, so that the VSD loss can provide more consistent gradient guidance conditioned on $t_{s}$. Specifically, we feed the randomly sampled $t_{s}$ and the LQ image into the student model to obtain $\hat{z}=G_{\theta}(x_{L},t_{s})$. Then, $t_{s}$ is mapped to $t_{v}$ by:
$t_{v}=\lambda t_{s}+\gamma,t_{s}\in[0,999],$ (4)
where the $\lambda$ and $\gamma$ are the hyperparameters.
We then add noise to $\hat{z}$ corresponding to $t_{v}$ to obtain $\hat{z}_{t_{v}}=\alpha_{t_{v}}\hat{z}+\beta_{t_{v}}\epsilon$, which is then passed through the teacher model and the LoRA model with $t_{v}$ to compute the TAVSD loss:
$\begin{gathered}\nabla_{\theta}\mathcal{L}_{TAVSD}(\hat{z},c,t_{v})=\mathbb{E}_{\epsilon}[\omega(t_{v})(\epsilon_{\psi}(\hat{z}_{t_{v}};t_{v},c)\\ -\epsilon_{\phi}(\hat{z}_{t_{v}};t_{v},c))\frac{\partial\hat{z}}{\partial\theta}].\end{gathered}$ (5)
By leveraging the TAVSD loss, the model can naturally balance generation and fidelity in the Real-ISR task simply by varying the timestep condition $t_{s}$ input to the model.
Figure 5: Visual comparisons between our method and other Real-ISR methods. Please zoom in for a better view.
Training Loss
We train the student model with reconstruction and regression losses. To avoid gradient inconsistency arising from the ill-posed problem of the Real-ISR task (Liang, Zeng, and Zhang 2022) while fully leveraging the teacher model knowledge introduced by VASVD, we first apply a Gaussian blur to both the reconstructed image and the HQ image $x_{H}$ before computing the MSE loss. This ensures that the $x_{H}$ only supervises the low-frequency content of the reconstruction, helping to preserve high-frequency details. We adopt a larger blur kernel for larger timesteps $t_{s}$, which enhances the trade-off between fidelity and generation:
$\mathcal{L}_{MSE}^{blur}=\mathcal{L}_{MSE}\left(G_{\theta}(x_{L})*G_{t_{s}},x_{H}*G_{t_{s}}\right).$ (6)
Where * denotes the convolution operation, $G_{t_{s}}$ is the convolution kernel whose size is determined by $t_{s}$. This loss and the LPIPS loss, forms the reconstruction loss:
$\mathcal{L}_{Rec}=\mathcal{L}_{MSE}^{blur}+\mathcal{L}_{LPIPS}(G_{\theta}(x_{L}),x_{H}).$ (7)
For the regression loss, we adopt the TAVSD loss in Eq. 5 to improve the realism of the generated results. The overall loss for the student model is:
$\mathcal{L}_{Stu}=\mathcal{L}_{Rec}+\lambda_{TAVSD}\cdot\mathcal{L}_{TAVSD}.$ (8)
For the LoRA model, we adopt the original diffusion loss:
$\mathcal{L}_{Diff}(\hat{\mathbf{z}},c_{y})=\mathbb{E}_{t,\epsilon}\left[\left\|\epsilon_{\phi}(\hat{\mathbf{z}}_{t};t,c_{y})-\epsilon^{\prime}\right\|^{2}\right],$ (9)
where $\epsilon^{\prime}$ is the Gaussian noise as the training target for the denoising network.
Experiments
Experimental Setup
Training. We use LSDIR (Li et al. 2023) as the training data with the $512\times 512$ patch size. To generate paired HQ-LQ training data, we follow the degradation pipeline from Real-ESRGAN (Wang et al. 2021). We use AdamW optimizer (Loshchilov and Hutter 2017) with a learning rate $5\times 10^{-5}$ and set LoRA rank to 4 for both the student model and LoRA model. We employ the SD 2.1-base as the pre-trained diffusion model and fine-tune it for 2k iterations using 8 NVIDIA A40 GPUs with a batch size of 24.
Test Dataset. We evaluate our method in both synthetic and real-world dataset. For the synthetic dataset, we randomly crop 3K patches with a resolution of 512 × 512 from the DIV2K (Agustsson and Timofte 2017) validation set and synthesize LQ data using the same pipeline as that in training. For real-world data, we employ RealSR (Cai et al. 2019), DrealSR (Wei et al. 2020), and RealLR200 (Wu et al. 2024c). We center-crop RealSR (Cai et al. 2019) and DrealSR (Wei et al. 2020) datasets with size $128\times 128$ for LQ images and $512\times 512$ for HQ image. For RealLR200 (Wu et al. 2024c) dataset, since the corresponding HQ images are unavailable, we perform only a $128\times 128$ center-crop on the LQ images.
Evaluation Metrics. We utilize several reference and non-reference metrics to evaluate the performance of various methods on the test data. For the reference measures, we employ PSNR, SSIM (Wang et al. 2004), and LPIPS (Zhang et al. 2018a) to measure image fidelity. For the non-reference measures, we employ CLIPIQA (Wang, Chan, and Loy 2023), MUSIQ ((Ke et al. 2021), MANIQA (Yang et al. 2022), TOPIQ (Chen et al. 2024), and QALIGN (Wu et al. 2024a) to measure image quality.
Compared Methods. We compare our method with several multi-step diffusion-based methods StableSR (Wang et al. 2023a), DiffBIR (Lin et al. 2023), SeeSR (Wu et al. 2024c)), and one-step diffusion-based methods SinSR (Wang et al. 2024), OSEDiff (Wu et al. 2024b), S3Diff (Zhang et al. 2024), AdcSR (Chen et al. 2025), TSDSR (Dong et al. 2025), and PisaSR (Sun et al. 2025). All comparative results are obtained using publicly released code for testing.
Comparisons with State-of-the-art Methods
Quantitative Comparisons. We set up the timestep condition $t_{s}=500$ in our method, and show the quantitative comparisons on the four synthetic and real-world datasets in Tab. 1. We have the following observations: (1) TADSR achieves the highest no-reference scores across four datasets, except for the MUSIQ on DIV2K-Val. This demonstrates that TADSR can more effectively leverage the generative priors from SD to produce more realistic results. Notably, TADSR is the only one-step methods that consistently outperforms multi-step methods on all no-reference metrics, achieving both efficiency and perceptual quality. (2) TADSR maintains PSNR values comparable to other SD-based one-step methods, indicating a good balance between fidelity and realism. (3) TADSR shows clear improvements over other SD-based one-step methods on CLIPIQA and TOPIQ, highlighting its superior semantic awareness and generative capability.
Figure 6: (a) Quantitative metrics of our method under different timestep condition $t_{s}$, evaluated on the DrealSR dataset. (b) Comparison of our method under different timesteps $t_{s}$, PisaSR under different semantic guidance weights $\lambda_{sem}$, and other one-step diffusion-based Real-ISR methods, evaluated on the DrealSR dataset.
Qualitative Comparisons. Fig. 5 shows the visual comparisons between our method and the other state-of-the-art Real-ISR methods. As shown in the first row, TADSR generates significantly more natural and sharper textures from heavily degraded LQ images, especially in facial regions such as the teeth, eyes, and eyebrows, demonstrating its strong semantic generation capability. In the second row, the digits and letters produced by TADSR appear much clearer, showcasing its superior degradation removal ability while preserving fidelity. In the third row, TADSR yields more natural results around the eagle’s eyes and beak. In the fourth row, only TADSR accurately restored natural-looking facial features such as the nose, mouth, and chin. Other methods generally suffered from degradation, resulting in some distortion, and failed to reconstruct a plausible chin structure. Overall, thanks to the ability to distill generative priors from SD more effectively in TAVSD loss, TADSR can produce natural and realistic results in a single diffusion step. Compared to other methods, it achieves strong perceptual quality while maintaining high efficiency.
Ablation Study
Impact of Different Timestep Condition. As shown in Fig. 6(a), we analyze the impact of timestep $t_{s}$ in our method on both reference and no-reference metrics. As $t_{s}$ increases, PSNR exhibits a decreasing trend while QALIGN shows an upward trend, indicating a trade-off where fidelity is sacrificed to enhance realism. This trade-off between fidelity and realism aligns with the function of $t_{s}$, as a larger $t_{s}$ means that TAVSD provides stronger generative guidance, while a smaller $t_{s}$ provides more fidelity-preserving guidance. Similar visual results can be observed in supplementary materials. Furthermore, we compare the results of our method under different $t_{s}$, PisaSR under different $\lambda_{sem}$ settings, and other one-step Real-ISR methods, as shown in Fig. 6(b). It can be observed that our method consistently lies in the top-right corner across different $t_{s}$. When $t_{s}$ equals 200, our method achieves 26.61dB PSNR, which is more than 1dB higher than SinSR, and QALIGN is significantly higher than SinSR. In contrast, although PisaSR can also achieve a PSNR of 29.60dB by tuning the $\lambda_{pix}=1.0$ and $\lambda_{sem}=0.6$, its QALIGN is only 2.91, which is similar to SinSR. This indicates that our method achieves a substantial improvement in fidelity with only a minimal compromise in realism.
Figure 7: Vision Comparisons of the ablation study on TAE and TAVSD. Baseline use the original VAE encoder in SD and VSD loss.
[…]
RealSR Baseline 24.39 70.22 0.6751 0.6391
w/o TAE 24.89 70.08 0.6857 0.6466
w/o TAVSD 24.84 70.96 0.6930 0.6553
Full 25.16 71.18 0.7283 0.7082
DrealSR Baseline 27.45 65.90 0.6887 0.6275
w/o TAE 27.95 65.95 0.7030 0.6396
w/o TAVSD 28.03 66.95 0.7015 0.6373
Full 28.39 67.02 0.7398 0.6758
Table 2: Quantitative Comparison of ablation study on TAVSD and TAE. Baseline uses the original VAE encoder in SD and VSD loss.
Impact of TAVSD and TAE. To validate the effectiveness of TAVSD and TAE, we conducted ablation studies by removing them. We employ the original VAE encoder in SD and the VSD loss as our baseline and conduct ablation studies by separately removing TAE and TAVSD. We use PSNR to evaluate fidelity and MANIQA, MUSIQ, and TOPIQ to assess realism. As shown in Tab. 2, we have the following three key observations: (1) After removing TAE, both reference and no-reference metrics decline, demonstrating that adapting the latent feature according to timesteps helps more fully utilize the generative priors in SD. (2) When TAVSD is ablated, all metrics similarly decrease, indicating that more consistent guidance from the teacher model better activates generative priors across different timesteps. (3) Baseline shows significant degradation in PSNR and moderate decline in others, proving that both TAE and TAVSD improve fidelity and realism. Additionally, Fig. 7 presents a visual comparison of our ablation studies, showing that both the absence of TAE/TAVSD leads to unrealistic parrot reconstructions while the baseline even produces visible artifacts. In contrast, our method produces realistic and natural results by fully exploiting the generative priors in SD.
Conclusion
In this paper, we propose TADSR, a one-step SD-based Real-ISR method. TADSR introduces a variable timestep $t_{s}$ into the student model and uses a Time-Aware VAE Encoder to fully utilize the generative priors in SD at different timesteps. To further distill the priors at different timesteps in SD to achieve varied SR effects, TADSR leverages the Time-Aware Variational Score Distillation to enable the teacher model to provide more consistent generative guidance condition on $t_{s}$. As a result, TADSR fully leverages the generative priors in SD and naturally achieves a controllable trade-off between fidelity and realism condition on $t_{s}$. Our experiments demonstrate that TADSR achieves state-of-the-art performance among all Real-ISR methods.